A woman has a ladder that is 13 feet long. If she sets the base of the ladder on level ground 5 feet from the side of a house, how many feet above the ground will the top of the ladder be when it rests against the house?
Answer by mananth(16949)

ladder = 13ft ===> hypotenuse
distance from wall = leg1===> 5 ft
height it touches the wall leg2=?
side2^2=hyp^2-side1^2
13^2-5^2=leg2 ^2
169-25= leg2^2
leg2=12 ft